A flat coil enclosing an area of 0.07m2 is rotating at 65 rev/s, with its axis of rotation perpendicular to a 0.23 T magnetic field.
(a) If there are 996 turns on the coil, what is the maximum voltage induced in the coil?
V

(b) When the maximum induced voltage occurs, what is the orientation of the coil with respect to the magnetic field?
The plane of the coil is perpendicular to the magnetic field.
The plane of the coil is parallel to the magnetic field.
The plane of the coil is at a 65° angle to the magnetic field.
The plane of the coil is at a 43° angle to the magnetic field.
Answer #1
area A = 0.07 m2 , ω= 65 rev/s = 65*2π rad/s, B =0.23 T, turns N = 996
a) maximum voltage = NωBA = 6.25*103 V
b) The plane of the coil is parallel tothe magnetic field.